# Agent Design Pattern Catalogue: A Collection of Architectural Patterns for Foundation Model based Agents

### v3

@@ 4.14 Debate-based Cooperation
Solution: Fig. 14 depicts a high-level graphical representation of debate-based cooperation. A user can send queries to an agent, which will then share the questions with other agents. Given the shared question, each agent generates its own initial responses, and subsequently, a round of debate will start between the agents. Agents will propagate their initial response in a decentralised manner to each other for verification, while also providing instructions and potential planning directions to construct a more comprehensive response based on inclusive and collective outcomes. In addition, agents may utilise a shared memory in certain circumstances, or allow each other to access the respective memory facilitating the debate. This debate process can be iterative to enhance the performance of all participating agents. Debate-based cooperation can end according to a predefined number of debate rounds, or the agents will continue the procedure until a consensus answer is obtained.
@@ 4.15 Multimodal Guardrails
Solution: Fig. 15 presents a simplified graphical representation of multimodal guardrails. Guardrails can be applied as an intermediate layer between the foundation model and all other components in a compound AI system. When users send prompts or other components (e.g. memory) transfer any message to the foundation mode, guardrails can first verify whether the information meets specific predefined requirements, only valid information will be delivered to the foundation model. For instance, personally identifiable information should be treated with care or removed to protect privacy. Guardrails can evaluate the contents either relying on predefined examples, or in a “reference-free” manner. Equivalently, when the foundation model creates results, the guardrails need to ensure that the responses do not include biased or irrespective information to users, or fulfil the particular requirements of other system components. Please note that a set of guardrails can be implemented where each of them is responsible for specialised interactions, e.g., information retrieval from datastore, validation of users’ input, external API invocation, etc. Meanwhile, guardrails are capable of processing multimodal data such as text, audio, video to provide comprehensive monitoring and control.
@@ 4.16 Tool/Agent Registry
Solution: Fig. 16 depicts how can an agent search external agents and tools via a tool/agent registry. A user prompts goals to an agent, which then decomposes the goals into fine-grained tasks. The agent can query the tool/agent registry, which is the main entry point for collecting and categorising various tools and agents regarding a series of metrics (e.g., capability, price, context window). Based on the returned information, the agent can employ and assign the tasks to respective tools and agents. Please note that a registry can be implemented in different manners, for instance, a coordinator agent with specific knowledge base, blockchain-based smart contract, etc., and a registry can be extended into a marketplace for tool/agent service trading.
@@ 4.17 Agent Adapter
Solution: Fig. 17 demonstrates a simplified graphical representation of agent adapter. Given user’s instructions, the agent generates a plan consisting of a set of tasks to achieve the user’s goals. In particular, the agent may employ diverse external tools to complete different tasks. However, tools have respective interfaces, which can be of different abstraction levels for the agent to deal with, or have specific format requirements, etc. Agent adapter can help invoke and manage these interfaces by converting the agent messages into required format or content, and vice versa. In particular, the adapter can retrieve tool manual or tutorial from datastore, to acquire available interfaces. It then transforms the agent outputs based on the interface requirements, and invokes the service [68]. Please note that fine-grained interface description can help agent to understand and hence improve the performance. The adapter also receives execution results from tools, which will be sent to the underlying foundation model for further analysis (e.g. task assignment to other tools, self-reflection for tool employment). For instance, the adapter can translate tasks into system messages when interacting with local file system, or capture and operate graphical user interface when playing a video game.
@@ 4.18 Agent Evaluator
Solution: Fig. 18 presents a simplified graphical representation of agent evaluator. Developers can deploy evaluator to assess the agent regarding responses and reasoning process at both design-time and runtime. Specifically, developers need to build up the evaluation pipeline, for instance, by defining specific scenario-based requirements, metrics and expected outputs from agents. Given particular context, the agent evaluator prepares context-specific test cases test cases (either searching from external resources or generating by itself), and performs evaluation on the agent components respectively. The evaluation results provide valuable feedback such as boundary cases, near-misses, etc., while developers can further fine-tune the agent or employ corresponding risk mitigation solutions, and also upgrade the evaluator based on the results.
@@
5 Discussion
This section discusses several insights based on our experiences during the pattern collection process and previous research projects.
@@ 5 Discussion
Integration with extant patterns. Integrating different patterns can help structure and develop comprehensive and trustworthy agents. In particular, the proposed pattern catalogue can be applied together with responsible AI patterns [70] to ensure the agents behave in a responsible manner. For instance, bill of materials registry can record the procurement of components such as guardrails, prompt/response optimiser, etc., offering a complete supply chain of foundation model-based agent development. Whilst, black box can be implemented to collect the inputs and outputs of foundation model-based agents at runtime. If any abnormal behaviour is detected or an audit is needed, the recorded information can provide evidence for the accountability process. In addition, a set of voting mechanisms can be leveraged via blockchain smart contracts [71] for voting-based cooperation, to ensure transparent and secure interactions between agents.
@@
Compliance with regulations and standards. Preserving the alignment of agents with both international and domestic regulations and standards should be noted as a fundamental factor for developers to provide agent services in different countries and regions. European Parliament has approved the Artificial Intelligence Act [^29] https://artificialintelligenceact.eu/ which focuses on three risk categories of AI applications. NIST released a draft publication for managing the risks of generative AI [^30] https://www.nist.gov/itl/ai-risk-management-framework, and ISO published an international standard for implementing and maintaining AI management systems [72]. Future work can extract and analyse the requirements within each regulation and standard, and map the proposed pattern catalogue to the requirements.
@@
Evaluation of foundation model-based agents. Evaluations on agents and the underlying foundation models are significant to ensure they behave as intended. The majority of pattern benefits and drawbacks are software quality attributes, which still require quantification for fine-grained metrics and rubrics. For instance, accountability can be further divided into three criteria of responsibility, auditability, and redressability, and each criterion has its own process, resource, and product metrics [73]. Proper quantification can promote the evaluation of agents and validate the effectiveness of applied patterns. Please note that the evaluation of foundation model-based agents can be categorised into system-level and component-level [74], while each granularity has respective testing methods and benchmarking.

### v4

@@ 4.14 Debate-based Cooperation
Solution: Fig. 14 depicts a high-level graphical representation of debate-based cooperation. A user can send queries to an agent, which will then share the questions with other agents. Given the shared question, each agent generates its own initial responses, and subsequently, a round of debate will start between the agents. Agents will propagate their initial response in a decentralised manner to each other for verification, while also providing instructions and potential directions to construct a more comprehensive response based on inclusive and collective outcomes. In addition, agents may utilise a shared memory in certain circumstances, or allow each other to access the respective memory for facilitating the debate. This debate process can be iterative to enhance the performance of all participating agents. Debate-based cooperation can end according to a predefined number of debate rounds, or the agents will continue the procedure until a consensus answer is obtained.
@@
  • Li et al. [64] propose a framework including peer rank and discussion between agents, to mitigate the biases in automated evaluation process.
@@ 4.15 Multimodal Guardrails
Solution: Fig. 15 presents a simplified graphical representation of multimodal guardrails. Guardrails can be applied as an intermediate layer between the foundation model and all other components in a compound AI system. When users input prompts or other components (e.g. memory) send any message to the foundation model, guardrails can first verify whether the information meets specific predefined requirements. Only valid information are delivered to the foundation model, while risky or sensitive data will be processed before being transferred. For instance, personally identifiable information should be treated with care or removed to protect privacy. Guardrails can evaluate the contents either relying on predefined examples, or in a “reference-free” manner. Equivalently, when the foundation model creates results, guardrails need to ensure that the responses do not include biased or irrespective information to users, or fulfil the particular requirements of other system components. Please note that a set of guardrails can be implemented where each of them is responsible for specialised interactions, e.g., information retrieval from datastore, validation of users’ input, external API invocation, etc. Meanwhile, guardrails are capable of processing multimodal data such as text, audio, video to provide comprehensive monitoring and control.
@@ 4.16 Tool/Agent Registry
Solution: Fig. 16 depicts how an agent searches external agents and tools via a tool/agent registry. A user prompts goals to an agent, which then decomposes the goals into fine-grained tasks. The agent can query the tool/agent registry, which is the main entry point for collecting and categorising various tools and agents regarding a series of metrics (e.g., domain-specific capability, price, context window). Based on the returned information, the agent can employ and assign the tasks to respective tools and agents. Please note that a registry can be implemented in different manners, for instance, a coordinator agent with specific knowledge base, blockchain-based smart contract, etc., and a registry can be extended into a marketplace for tool/agent service trading.
@@ 4.17 Agent Adapter
Solution: Fig. 17 demonstrates a simplified graphical representation of agent adapter. Given user’s instructions, the agent generates a plan consisting of a set of tasks to achieve the user’s goals. In particular, the agent may employ diverse external tools to complete different tasks. However, tools have respective interfaces, which can be of different abstraction levels for the agent to deal with, or have specific format requirements, etc. Agent adapter can help invoke and manage these interfaces by converting the agent messages into required format or content, and vice versa. In particular, the adapter can retrieve tool manual or tutorial from datastore, to acquire available interfaces and learn the usage. It then transforms the agent outputs based on the interface requirements and invokes the service [70]. Please note that fine-grained interface description can enhance agent understanding and hence improve the performance. The adapter also receives execution results from tools, which will be sent to the underlying foundation model for further analysis (e.g. task assignment to other tools, self-reflection for tool employment). For instance, the adapter can translate tasks into system messages when interacting with local file system, or capture and operate graphical user interface when playing a video game.
@@ 4.18 Agent Evaluator
Solution: Fig. 18 presents a simplified graphical representation of agent evaluator. Developers can deploy evaluator to assess the agent regarding responses and reasoning process at both design-time and runtime. Specifically, developers need to build up the evaluation pipeline, for instance, by defining specific scenario-based requirements, metrics and expected outputs from agents. Given particular context, the agent evaluator prepares context-specific test cases (either searching from external resources or generating by itself), and performs evaluation on the agent components respectively. The evaluation results provide valuable feedback such as boundary cases, near-misses, etc., while developers can further fine-tune the agent or employ corresponding risk mitigation solutions, and also upgrade the evaluator based on the results.
@@
5 Lessons Learned From the Pattern Catalogue
In this section, we first propose a decision model for selecting the 18 identified patterns to enhance the development of foundation mode based agents. Afterwards, we would like to share our experiences on the application of certain patterns in research projects as case studies. Finally, we discuss several insights we learned during the pattern collection process and previous research works.
Figure 19: Decision model for agent design pattern catalogue.
5.1 Decision Model
Fig. 19 illustrates a decision model which visualises the selection process of different patterns for practitioners. In particular, for a design problem, each decision can map to the corresponding solution space, which is regarded as the alternative pattern(s) for those in the opposite decision’s solution space. Further, within the same solution space, there can be multiple patterns serving as complements of each other. The strengths and trade-offs of each pattern are highlighted in green and orange colour respectively. Please note that the decision model omits the common attributes for alternative patterns as either decision will incorporate these attributes as consequences, and it centralises the shared attributes for complementary patterns for brevity. In addition, there is no specific selection or application sequence, considering the patterns aim to facilitate the design and development of different architectural components, which can be implemented via a decoupling strategy. A brief explanation is provided as follows.
If the agent is expected to capture users’ environmental information as supplementary context, proactive goal creator can be applied to analyse users’ goals based on the data captured by a series of sensors. In this case, the agent can also serve users with specified disabilities. Whilst, passive goal creator can provide a simple and efficient dialogue interface to interact with users. Prompt/response optimiser can enhance the goal alignment by refining users’ instructions into standardised prompts. Meanwhile, the agent can retrieve more information from external knowledge base via retrieval augmented generation whenever a component requires additional context. Besides, a component can query the incorporated foundation model for a single purpose multiple times (incremental model querying) based on either user-specified requirements or system configuration, to provide supplementary context and hence improving model reasoning certainty, or just query once (one-shot model querying) due to limited budget for model calling.
For a particular goal, agents can create a linear plan via single-path plan generator for efficiency, or a complex plan in which each step has different options via multi-path plan generator to ensure inclusiveness and alignment to human preferences. A generated plan can be assessed via multiple solutions (i.e. self-reflection, cross-reflection and human reflection) for ensuring plan correctness and feasibility, and improving agent reasoning certainty and explainability, while each solution has respective strengths and trade-offs. Further, multiple agents can be employed for reflection, and they can interact with each other in terms of voting-based, role-based and debate-based cooperation schemes.
When the underlying foundation model is queried, malicious inputs can affect the reasoning process, and the model may continuously learn and generate skewed outputs. Multimodal guardrails can provide a layer between the foundation model and other components by inspecting the model inputs and outputs through both rule-based and AI-based examination. In case that external tools or agents are leveraged for certain tasks, a tool/agent registry can enhance discoverability, and ensure the efficiency and appropriateness of tools. Agent adapter guarantees the interoperability and adaptability of an agent to call external services whilst reducing the cost for manual development and maintenance. Finally, agent evaluator can be utilised to assess the agent’s functional suitability before release.
5.2 Pattern Application
In this section, we would like to discuss and share our experiences of applying the identified patterns in research projects. Researchers joined a project in which we are designing and developing an agent platform for creating and operating bespoke foundation model based agents with continuous learning capability. Specifically, the platform is divided into multiple products and components where different patterns can be applied.
Universal Task Assistant [^29] https://apputa.online/, developed by Xie et al., can help users discover and learn the usage of different mobile apps. This product applied proactive goal creator to capture the mobile UI screen and detect the included elements, then it can perform actions to complete the user’s tasks. Cheng et al. proposed an AI chain integrated development environment, Prompt Sapper [72], for practitioners to properly and seamlessly develop their own FM-based AI chain services. In particular, the Prompt Sapper co-pilot implemented passive goal creator for eliciting users’ requirements, prompt/response optimiser for refining users’ task descriptions and generating AI chain skeleton, and tool/agent registry for managing all available artifacts. Shamsujjoha et al. [73] further explored multimodal guardrails by devising a taxonomy of guardrails, which provides comparative analysis of diverse design options in terms of the actions, targets, scopes, rules, autonomy, modalities, and underlying techniques when incorporating guardrails in foundation model based systems. Whilst, Xia et al. [74] demonstrated an AI system evaluation framework, categorising the evaluation of foundation model based applications into system-level and component-level, each granularity has respective testing methods and benchmarking. The framework can offer guidance for the application of agent evaluator.
5.3 Discussion
@@
Integration with extant patterns. Integrating different patterns can help structure and develop comprehensive and trustworthy agents. In particular, the proposed pattern catalogue can be applied together with responsible AI patterns [75] to ensure the agents behave in a responsible manner. For example, bill of materials registry can record the procurement of components such as guardrails, prompt/response optimiser, etc., offering a complete supply chain of foundation model-based agent development. Whilst, black box can be implemented to collect the inputs and outputs of foundation model-based agents at runtime. If any abnormal behaviour is detected or an audit is needed, the recorded information can provide evidence for the accountability process. Further, the collaboration patterns among agents can refer to existing social learning evolution models [76]. For instance, selecting suitable topologies and implementing effective control mechanisms can enhance the design of multi-agent workflows. Agents might also benefit from imitative learning in debate-based cooperation, enabling dynamic adaptation and knowledge transfer. In addition, a set of voting mechanisms can be leveraged via blockchain smart contracts [77] for voting-based cooperation, to ensure transparent and secure interactions between agents.
@@
Compliance with regulations and standards. Preserving the alignment of agents with both international and domestic regulations and standards should be noted as a fundamental factor for developers to provide agent services in different countries and regions. European Parliament has approved the Artificial Intelligence Act [^30] https://artificialintelligenceact.eu/ which focuses on four risk categories of AI applications. NIST released a draft publication for managing the risks of generative AI [^31] https://www.nist.gov/itl/ai-risk-management-framework, and ISO published an international standard for implementing and maintaining AI management systems [78]. Future work can extract and analyse the requirements within each regulation and standard, and map the proposed pattern catalogue to the requirements. In particular, we are adopting the concept of DevOps into FM-based agents, where both agent creation and operation information will be recorded for further auditing. Meanwhile, the agent workflow generation also requires inspection that whether it adheres to domain specifications.
@@
Evaluation of foundation model-based agents. Evaluations on agents and the underlying foundation models are significant to ensure they behave as intended. The majority of pattern benefits and drawbacks are software quality attributes, which still require quantification for fine-grained metrics and rubrics. For instance, accountability can be further divided into three criteria of responsibility, auditability, and redressability, and each criterion has its own process, resource, and product metrics [79]. Proper quantification can promote the evaluation of agents and validate the effectiveness of applied patterns. Further, authors are also exploring real-time exception identification and handling in agent workflow execution, which includes rationalising task dependencies, external tool API version matching, etc.